OpenStudy (anonymous):

find the circumference of a circle who's area is 400pie^2

OpenStudy (anonymous):

Area = pir^2 so: 400pie^2 =pie r^2 400pie=r^2 r=sqrt 400pie

OpenStudy (anonymous):

circumference = pi*d=pi*2r=pi"2sqrt 400pie